OpenStudy (omgthatgirl):

Find the first three output values of the fractal generating function f(z)=z^2+2+i a. 2+i, 2+5i, -2595+205i b. 2+i, 5+5i, -2595+205i c. 2+i, 5+5i, 2+51i d. 5+5i, 2+51i, -2595+205i

OpenStudy (perl):

is the 'seed' value z = 0 ?

OpenStudy (omgthatgirl):

i thnk so

OpenStudy (perl):

ok plug in z = 0 into f(z)

OpenStudy (omgthatgirl):

k

OpenStudy (omgthatgirl):

i think its d

OpenStudy (perl):

f(0) = 0^2 + 2 + i = 2 + i so the first value is 2 + i f(2+i) = (2+i)^2 + 2 + i = (2 + 2i)(2+2i) + 2 + i = 5 + 5i

OpenStudy (perl):

so the next ouput is 5 + 5i

OpenStudy (omgthatgirl):

wourl the third be -2595?